OpenStudy (akohl):

@abhisar A 76-kg ball is tied to one end of a massless string of length The other end of the string is tied to a nail in the center of a table. The ball is set into uniform circular motion around the nail, and the tension in the string is 51 N. How long does the ball take to make one complete revolution? Assume friction with the table is negligible. 8.7 s 8.1 s 9.5 s 7.0 s

OpenStudy (akohl):

Not sure what to do

OpenStudy (abhisar):

radius not given ?

OpenStudy (akohl):

String length 1.3m

OpenStudy (abhisar):

aahh man..u made me mad !

OpenStudy (akohl):

Lol sorry

OpenStudy (abhisar):

The tension in the string will be equal to the centripetal force, \(\large\sf \frac{MV^2}{r}=51\)

OpenStudy (abhisar):

now find the value of V

OpenStudy (abhisar):

Then use \(\large v= \frac {2 \pi r } {T}\) to find T

OpenStudy (akohl):

I got 8.85 which isn't an option

OpenStudy (abhisar):

I am getting 8.7 !

OpenStudy (akohl):

Maybe I rounded a little that probably changed it thanks

OpenStudy (abhisar):